# Learning Diffusion Policies for Robotic Manipulation of Timber Joinery under Fabrication Uncertainty

### v1

@@ 3.1 Phase 1: Static experiments
The objective of the Phase 1 experiments is to show the application and performance of diffusion policies in a building-scale robotic assembly setting. For this experiment, we created a dataset of 100 demonstrations following the data collection method described in Section 2.2, with the exception that no demonstrations included error recovery cases and demonstrations were not post-processed. To simplify the assembly task, the mortise position was held fixed across all demonstrations and during rollouts. Through repeated experimentation, we identified the best-performing set of parameters (reported in Section 4.1), which were then used to initialize parameter tuning in the Phase 2 experiments described in the following subsection.
@@ 3.2 Phase 2: Uncertainty experiments
The objective of the Phase 2 experiments is to demonstrate the robustness of diffusion policies in handling fabrication uncertainties for contact-rich timber joinery. More specifically, we evaluate how the policies perform when the mortise is subjected to random position offsets simulating fabrication tolerances.
We created a dataset of 400 demonstrations following the data collection method outlined in Section 2.2. To simulate fabrication tolerances, the position of the mortise in these demonstrations was perturbed by a displacement vector sampled in polar coordinates. The angular component was drawn uniformly from $0,2\pi)$ radians, and the radial component was drawn uniformly from $[0,10]$ mm. Unlike uniform sampling over the area of a disk, this approach yields a higher density of samples near the origin, biasing the distribution toward smaller offsets while still allowing for the possibility of larger deviations.
During rollouts, we evaluated each policy at three goal offset distances: 0 mm (no offset), 5 mm, and 10 mm. These values were selected to evaluate the policies under no uncertainty (0 mm), uncertainty within the demonstration distribution (5 mm), and uncertainty at the edge of the demonstration distribution (10 mm). For each offset condition, the goal position was randomized uniformly along the circumference of a circle with a radius of the specified distance. The success rate of a policy was computed separately for each offset distance, and its overall performance was reported as the average success rate across all three conditions.
In the first part of the Phase 2 experiments, we systematically tuned the parameters of the diffusion policy to maximize the success rate. Starting from a candidate configuration derived from the Phase 1 experiments (Section 3.1), we performed sequential, greedy parameter tuning, optimizing one parameter at a time while holding others fixed. The resulting best-performing policy is referred to as the full model. The impact of individual parameters on policy performance is reported in Section 4.2.1.
In the second part of the Phase 2 experiments, we conducted two ablation studies to assess the impact of F/T observations and demonstration count on policy performance. In the F/T ablation study, we evaluated the performance of the full model with F/T inputs masked (i.e., input F/T values are set to 0), as well as a variant trained using only pose observations. For the demonstration ablation, we trained policies using 25, 50, 100, and 200 demonstrations, while keeping all other parameters identical to the full model. The results of these studies are presented in Sections 4.2.2 and 4.2.3.
@@ 4 Results and discussion
Figure 8: Smoothed mean squared error (MSE) training and validation loss curves for Policy 4, averaged then smoothed across all training iterations. The shaded area represents $\pm 1$ standard deviation.
@@ 4.1 Phase 1: Static experiments
After repeated testing and parameter tuning, the policy achieved a 100% average success rate over 20 rollouts using the parameter values listed in Table 1. This performance demonstrates the capability of diffusion policies to reliably assemble a timber mortise and tenon joint under no uncertainty. These parameter values were subsequently used as the initialization point for parameter tuning in the Phase 2 experiments.
@@
During the first phase of the experiments, we also evaluated encoder networks for pose and F/T inputs to the diffusion policy. Each encoder consisted of a single fully connected layer that projected the raw input vector to a latent representation of fixed length (e.g., 64 or 128 units). The aim was to learn a compact feature embedding prior to policy integration. However, policies incorporating these encoders underperformed relative to those that directly concatenated the unprocessed pose and F/T observations. Future research may explore the impact of more sophisticated encoders, such as multi-layer, frequency-based, or attention-based architectures, on policy performance Lee et al. [2020], Hou et al. [2025].
@@
4.2 Phase 2: Uncertainty experiments
4.2.1 Full model parameter tuning
The sequential parameter tuning in Phase 2 was limited to three stages: (1) pose and F/T observation horizons; (2) action prediction horizon; (3) number of inference steps and DDIM/DDPM interpolation. The learning rate, weight decay, and forward noising steps were fixed at the values determined in Phase 1.
@@ 4.2.1 Full model parameter tuning
Table 3 summarizes the final stage of rollout parameter tuning, evaluating different combinations of inference steps and DDIM/DDPM interpolation values. These tests were conducted only for mortise offsets of 5 mm. The initial parameter values selected from the Phase 1 experiments achieved the highest performance, with a 65% average success rate, substantially outperforming other tested configurations.
Table 3: Success rates for rollout parameter tuning.
@@
4.2.2 Force/torque ablation study
Table 4 summarizes the results of the F/T ablation study, and Fig. 10 illustrates the contribution of F/T feedback to task performance. As expected, both ablated policies underperformed relative to the full model. However, a notable pattern emerged when comparing the pose-only policy to the masked full model. When F/T inputs were masked in the full model, the tenon consistently collided with the mortise, reflecting the model’s inability to detect contact without force feedback. Surprisingly, the pose-only policy occasionally succeeded under the 10 mm mortise offset condition, despite lacking force information. The large standard error in model variance (Fig. 10) for this condition suggests that these successes were likely the result of random factors rather than reliable inference. This highlights the critical role of force feedback in consistent policy performance, particularly under larger fabrication uncertainties.
@@ 4.2.2 Force/torque ablation study
  • *The forces were masked during rollouts.
@@
4.2.3 Demonstration quantity ablation study
Table 5 summarizes the results of the demonstration quantity ablation study. Policies trained on only 25 demonstrations failed to fully converge and achieved no success during rollouts, and were therefore excluded from the reported results. Fig. 11 shows a pronounced drop in success rate at the 10 mm mortise offset when the number of demonstrations was reduced, likely due to insufficient coverage of the state-action space in the demonstration data. The results suggest a performance inflection point between 200 and 400 demonstrations, beyond which the policy gains a substantial improvement in its ability to handle positional uncertainty. This trend indicates that under higher uncertainty, diffusion policies may require a critical mass of demonstrations to adequately represent recovery behaviors.
Notably, the total number of demonstrations required to achieve robust performance was much higher than anticipated. This is likely attributable to the trajectory subsampling strategy used during policy learning (Section 2.3), which significantly reduced the total number of training samples available to the policy. Future research should utilize alternative data preprocessing strategies to maintain a sufficient quantity of training data to enable reliable policy learning.
Table 5: Success rates for the demonstration quantity ablation study.
@@ 5 Conclusion
The experimental results presented in this work demonstrate that diffusion policies can achieve stable convergence and high success rates in a building-scale contact-rich assembly task representative of real-world construction. In Phase 1, we established a baseline by identifying parameter settings capable of achieving 100% average success rate in a deterministic mortise and tenon joint assembly task. In Phase 2, we evaluated policy robustness under controlled fabrication uncertainties, identifying parameter configurations that improved generalization to offsets of the mortise position, quantifying the critical role of force/torque feedback, and determining the demonstration quantity threshold necessary for reliable performance.
Together, these findings address the study’s core objectives: they validate the applicability of diffusion policies to building-scale robotic assembly and systematically characterize their robustness under fabrication uncertainty. Beyond the performative metrics, the experiments also yielded actionable insights for parameter tuning, sensor fusion, and demonstration count, forming a generalizable methodology for future contact-rich construction and assembly tasks. By demonstrating that diffusion policies can be systemically adapted and tuned to handle fabrication uncertainties without manual intervention, this research establishes a scalable, feedback-driven robot learning framework that addresses not only the technical demands of precision assembly, but also broader industry challenges such as labor shortages and the need to reduce reliance on physically taxing, repetitive, and high-skill manual work. Beyond timber joinery, the presented methodology can be generalized to a wide range of construction and manufacturing scenarios, from complex timber connections and pipe fitting to light metal framing, where expert craftsmanship has traditionally been essential. In doing so, this work advances robotic construction under uncertainty while contributing to safer and more efficient building practices, positioning robots as capable collaborators in the evolving construction workforce.
@@ 5.1 Limitations and future work
While this work demonstrates the feasibility and robustness of applying sensory–motor diffusion policy learning to contact-rich robotic assembly at construction scale, several limitations remain; our approach relies exclusively on robot pose and force data to train the policy and evaluate robustness, whereas diffusion policies for robot planning and control are commonly formulated as visuomotor policies using end-to-end mappings from RGB or tactile images to robot commands Chi et al. [2024a, b]. Our simplification of the sensory input space increases training speed but may limit the model’s ability to generalize to visually complex or occlusion-prone construction scenarios, such as joint assembly in cluttered framing environments or alignment tasks in multi-material panel installations. Moreover, incorporating expressive state-of-the-art learning models, such as diffusion transformers or vision-language-action (VLA) models Octo Model Team et al. [2024], Kim et al. [2024], could enhance representation learning, enabling better generalization, scalability, and transfer across tasks. These capabilities are particularly relevant for building-scale construction where robots must integrate multiple sensing modalities and operate across varied materials, such as combining timber, metal fasteners, and composite claddings in a single assembly sequence.
Furthermore, a second limitation stems from our reliance on single-task behavior cloning from human demonstrations, which introduces both practical and algorithmic challenges. Collecting high-quality demonstrations and conducting real-world rollouts is time-intensive and can cause material degradation, especially in contact-sensitive scenarios or fragile materials like softwoods, architectural veneers, or insulation panels. Moreover, the learned policy is inherently bounded by the quality and diversity of human performance, raising the question of how to scale beyond demonstrator capabilities. This is especially critical for more dexterous and varied AEC tasks, such as assembling shingle patterns on curved surfaces, handling flexible ductwork, or manipulating deformable architectural fabrics Craney and Adel [2020], Graser et al. [2021]. Future work could address these limitations by exploring hybrid learning paradigms that combine demonstration-based training with simulation pretraining, self-improvement cycles, or reinforcement learning, enabling policies to acquire physical reasoning skills and recover from unseen conditions TRI LBM Team et al. [2025]. Additionally, emerging research on general-purpose robot foundation models, such as visuomotor foundation models and Large Behavior Models (LBMs) TRI LBM Team et al. [2025], Zhao et al. [2023], presents a promising direction for learning transferable policies across a broad range of dexterous manipulation tasks.

### v2

@@ 4.3 Phase 1: Static experiments
The objective of the Phase 1 experiments is to evaluate the applicability and baseline performance of diffusion policies for contact-rich manipulation of the mortise and tenon joint. The policies were trained on the first dataset of 100 demonstrations (Section 4.2). Through repeated experimentation, we identified the best-performing parameter set (reported in Section 5.1) and used it to initialize parameter tuning in the Phase 2 experiments described in the following subsection.
@@ 4.4 Phase 2: Uncertainty experiments
The objective of the Phase 2 experiments is to evaluate the robustness of diffusion policies in handling uncertainties arising from positioning errors, tolerance accumulation, and material imperfections commonly encountered in construction-scale assembly. Specifically, we assess policy performance when the mortise is subjected to randomized positional offsets to simulate fabrication uncertainty. The mortise orientation is not randomized and remains constant. The policies for these experiments were trained using the second dataset of 400 demonstrations (Section 4.2).
During rollouts, we evaluated each policy with varying mortise position perturbations: 0 mm (no perturbation), 5 mm, and 10 mm. These values were selected to evaluate the policies under no uncertainty (0 mm), uncertainty within the demonstration distribution (5 mm), and uncertainty at the edge of the demonstration distribution (10 mm). For each offset condition, the mortise position was uniformly randomized along the circumference of a circle with a radius equal to the specified distance. The Avg. SR of a policy was computed separately for each offset distance, and its overall performance was reported as the average total success rate (Avg. total SR) across all three conditions.
In the first part of the Phase 2 experiments, we systematically tuned the diffusion policy parameters to maximize the success rate. Starting from a candidate configuration derived from the Phase 1 experiments (Section 4.3), we performed sequential, greedy parameter tuning, optimizing one parameter at a time while holding others fixed. The resulting best-performing policy is referred to as the full model. The impact of individual parameters on policy performance is reported in Section 5.2.1.
In the second part of the Phase 2 experiments, we conducted two studies to assess the impact of F/T observations and demonstration count on policy performance. In the F/T ablation study, we evaluated the performance of the full model with F/T inputs masked (i.e., F/T values set to 0) during inference, as well as a variant trained only on pose observations. For the demonstration count, we trained policies using 25, 50, 100, and 200 demonstrations (each including 25% error recovery cases as explained in Section 4.2), while keeping all other parameters identical to the full model. The results of these studies are presented in Sections 5.2.2 and 5.2.3.
@@ 5 Results and discussion
Figure 8: Smoothed mean squared error (MSE) training and validation loss curves for Policy 4, averaged and then smoothed across all training iterations. The shaded area represents $\pm 1$ standard deviation.
@@ 5.1 Phase 1: Static experiments
In Phase 1, after repeated testing and initial parameter tuning, the policy achieved a 100% average success rate over 20 rollouts using the parameter values listed in Table 1. This performance demonstrates the capability of diffusion policies to reliably assemble a timber mortise and tenon joint under no uncertainty. These parameter values were subsequently used as the initialization point for parameter tuning in the Phase 2 experiments.
@@
During the Phase 1 experiments, we also evaluated encoder networks for pose and F/T inputs to the diffusion policy. Each encoder consisted of a single fully connected layer that projected the raw input vector to a latent representation of fixed length (e.g., 64 or 128 units). The aim was to learn a compact feature embedding prior to policy integration. However, policies incorporating these encoders underperformed relative to those that directly concatenated the unprocessed pose and F/T observations. Future research may explore the impact of more sophisticated encoders, such as multi-layer, frequency-based, or attention-based architectures, on policy performance Lee et al. [2020], Hou et al. [2025].
@@
5.2 Phase 2: Uncertainty experiments
In Phase 2, we first perform systematic, sequential parameter tuning, starting from the Phase 1 configuration to identify the best-performing diffusion policy (referred to as the full model). Second, we analyze the impact of F/T observations and the number of demonstrations on policy performance.
5.2.1 Full model parameter tuning
The sequential parameter tuning during the Phase 2 experiments was limited to three stages: 1) pose and F/T observation horizons; 2) action prediction horizon; 3) number of inference steps and DDIM/DDPM interpolation. The learning rate, weight decay, and forward noising steps were fixed at the values determined during the Phase 1 experiments.
Table 3 summarizes the final stage of parameter tuning, evaluating different combinations of inference step $K_{inf}$ and DDIM/DDPM interpolation $\eta$. These tests were conducted only for mortise offsets of 5 mm, as the overall trend in success rates was consistent across other offsets. The initial parameter values selected from the Phase 1 experiments (i.e., $K_{inf}$ = 32 and $\eta$ = 0.5) achieved the highest performance, with a 65% average success rate, substantially outperforming other tested configurations.
Table 3: Success rates for inference parameter tuning.
@@ 5.2.1 Full model parameter tuning
5.2.2 Force/torque ablation
Table 4 summarizes the results of the F/T ablation study, and Fig. 10 illustrates the contribution of F/T feedback to task performance. As expected, both ablated policies underperformed relative to the full model. However, a notable pattern emerged when comparing the pose-only policy to the masked full model. When F/T inputs were masked in the full model, the tenon consistently collided with the mortise, reflecting the model’s inability to detect contact without force feedback. Surprisingly, the pose-only policy occasionally succeeded under the 10 mm mortise offset condition, despite lacking force information. The large standard error in model variance (Fig. 10) for this condition suggests that these successes were likely the result of random factors rather than reliable inference. This highlights the critical role of force feedback in consistent policy performance, particularly under larger perturbations.
@@ 5.2.2 Force/torque ablation
  • *The forces were masked during inference.
@@
5.2.3 Demonstration count
Table 5 summarizes the results of the demonstration quantity study. Policies trained on only 25 demonstrations failed to fully converge and achieved no success during rollouts, and were therefore excluded from the reported results. Fig. 11 shows a pronounced drop in success rate at the 10 mm mortise offset when the number of demonstrations was reduced, likely due to insufficient coverage of the state-action space in the demonstration data. The results suggest a performance inflection point between 200 and 400 demonstrations, beyond which the policy gains a substantial improvement in its ability to handle positional uncertainty. This trend indicates that under higher uncertainty, diffusion policies may require a critical mass of demonstrations to adequately represent recovery behaviors.
Notably, the total number of demonstrations required to achieve robust performance was much higher than anticipated. This is likely attributable to the trajectory subsampling strategy used during policy training (Section 3.3), which significantly reduced the total number of training samples available to the policy. Future research should utilize alternative data preprocessing strategies to maintain a sufficient quantity of training data to enable more reliable policy learning.
Table 5: Success rates for various demonstration counts.
@@ 6 Conclusion
The experimental results presented in this work demonstrate that diffusion policies can achieve stable convergence and high success rates in a contact-rich assembly task representative of real-world construction using industrial robotic arms. In the Phase 1 experiments, we established a baseline by identifying parameter settings capable of achieving 100% average success rate in a deterministic mortise and tenon joint assembly task. In the Phase 2 experiments, we evaluated policy robustness under controlled mortise position perturbations, identifying parameter configurations that improved generalization to offsets of the mortise position, quantifying the critical role of force/torque feedback, and determining the demonstration quantity threshold necessary for reliable performance.
These findings validate the applicability of diffusion policies to building-scale robotic assembly using industrial robots and systematically characterize their robustness under fabrication uncertainty. Beyond the performative metrics, the experiments also yielded actionable insights for parameter tuning, sensor fusion, and demonstration count, forming a generalizable methodology for future contact-rich construction and assembly tasks. By demonstrating that diffusion policies can be systemically adapted and tuned to handle fabrication uncertainties without manual intervention, this research establishes a scalable, feedback-driven robot learning framework that addresses not only the technical demands of precision assembly but also broader industry challenges such as labor shortages and the need to reduce reliance on physically taxing, repetitive, and high-skill manual work. Beyond timber joinery, the presented methodology can be generalized to a wide range of construction and manufacturing scenarios, from complex timber connections and pipe fitting to light metal framing, where expert craftsmanship has traditionally been essential. In doing so, this work advances robotic construction under uncertainty while contributing to safer and more efficient building practices, positioning robots as capable collaborators in the evolving construction workforce.
@@ 6.1 Limitations and future work
While this work demonstrated the feasibility and robustness of applying sensory–motor diffusion policy learning to contact-rich robotic assembly at construction scale, several limitations remain; our approach relies exclusively on robot pose and force data to train the policy and evaluate robustness, whereas diffusion policies for robot planning and control are commonly formulated as visuomotor policies using end-to-end mappings from RGB or tactile images to robot commands Chi et al. [2024a, b]. Our simplification of the sensory input space increases training speed but may limit the model’s ability to generalize to visually complex or occlusion-prone construction scenarios, such as joint assembly in cluttered framing environments or multi-alignment tasks. Moreover, incorporating expressive state-of-the-art learning models, such as diffusion transformers or vision-language-action models (VLAs) Physical Intelligence et al. [2025], Gemini Robotics Team et al. [2025], enables better generalization, scalability, and transfer across tasks. These capabilities are particularly relevant for building-scale construction where robots must integrate multiple sensing modalities and operate across varied materials, such as combining timber, metal fasteners, and composite claddings in a single assembly sequence.
Furthermore, a second limitation stems from our reliance on single-task behavior cloning from human demonstrations, which introduces both practical and algorithmic challenges. Collecting high-quality demonstrations and conducting real-world rollouts is time-intensive and can cause material degradation, especially in contact-sensitive scenarios or fragile materials like softwoods, architectural veneers, or insulation panels. Moreover, the learned policy is inherently bounded by the quality and diversity of human performance, raising the question of how to scale beyond demonstrator capabilities. This is especially critical for more dexterous and varied AEC tasks, such as assembling shingle patterns on curved surfaces, handling flexible ductwork, or manipulating deformable architectural fabrics Craney and Adel [2020], Graser et al. [2021]. Future work could address these limitations by exploring hybrid or hierarchical learning paradigms that combine demonstration-based training with simulation pretraining or reinforcement learning, enabling policies to acquire physical reasoning skills and recover from unseen conditions. Additionally, emerging research on general-purpose visuomotor foundation models such as VLAs or Large Behavior Models TRI LBM Team et al. [2025], Physical Intelligence et al. [2025] presents a promising direction for learning transferable policies across a broad range of dexterous manipulation tasks.